OpenStudy (micheal1234):

1over 4 a =3 a.0.75 b.3 c.12 d.16

OpenStudy (micheal1234):

one forth a=3

OpenStudy (jhannybean):

Is it \(\dfrac{1}{4}a = 3\) or \(\dfrac{1}{4a} = 3\) ?

OpenStudy (micheal1234):

the first one

OpenStudy (jhannybean):

Ok, I was wondering why the answer choices weren't matching up.

OpenStudy (jhannybean):

You want to isolate \(a\) onto one side of the equation, so you must multiply by the reciprocal. Multiply both sides of the equation by \(\dfrac{4}{1}\)

OpenStudy (micheal1234):

thanks :)

OpenStudy (micheal1234):

answer is 12